Franklin has $3.95 in nickels, dimes, and quarters. If he has three times as many dimes
as quarters, and five fewer nickels than dimes, how many coins of each type does he
have?

Answer by addingup(3677)

d = 3q
since d = 3q and nickels = dimes-3, we can say:
n = 3q-5
. - . - . - . - . - .
0.05n+0.10d+0.25q = 3.95 substitute per above:
0.05(3q-5)+0.10(3q)+0.25q = 3.95
0.15q-0.25+0.30q+0.25q = 3.95
group:
0.15q+0.30q+0.25q-0.25 = 3.95
0.7q = 4.20
q = 6 Franklin has 6 quarters
and
d = 3q = 3(6) = 18 he has 18 dimes
you do the nickels
:
John